0 0
Add a comment Improve this question Transcribed image text
Answer #1

6*10^-6F and 8.4*10^-6 F are in parallel then

equivalent capacitance is (6+8.4)\mu F= 14.4 \mu F

now the capacitor on right , left and equivalent capacitor are in series

total capacitance =\frac{1}{C_{t}}=\frac{1}{C_{1}}+\frac{1}{C_{2}}+\frac{1}{C_{3}}

C_{t}=\frac{14.4}{3}

C_{t}=4.8 \mu F

To find the charge we have to know the potential diffrence across the each capacitor.

as 6 \mu F and 8.4 \mu F are in parallel so voltage across them will be equal with equivalent capacitance.

therefore

V_{1}=\frac{C_{t}}{C_1}*V_{t}

V_{1}=\frac{4.8 \mu }{14.4 \mu }*90

V_{1}=30 V

as the capacitance is equal for other capacitor so the

voltage across them also equal with 30V

V2= 30 V

V3=30 V

therefore voltage across right , left 14.4 \mu F ,8.4\mu F and 6\mu F is 30 V

now charge across each capacitor as

on right and left 14.4 \mu F capacitor areQ=CV

Q14.4\mu F=14.4\mu F*30 V

Q_{14.4\mu F}=432 \mu C

Q_{8.4\mu F}=8.4\mu F*30 V

Q_{8.4\mu F}=252\mu C

Q_{6\mu F}= 6\mu F*30 V

Q_{6\mu F}=180\mu C